The length of a rectangle is 5m greater than the width.
The perimeter is 150m. find the width and length.
: The length of a rectangle is 5m greater than the width.
The perimeter is 150m. find the width and length.

Answer by orca(336)

Let x be the width of the rectangle, then the length of it is x + 5.
As the perimeter is 150, we can set up an equation using this fact:
2x + 2(x + 5) = 150
Solving for x, we obtain:
2x + 2x + 10 = 150
4x + 10 = 150
4x = 140
x = 35
so the width and length are 35 and (35 + 5 =)40 meters.